STIMULATED EMISSION PUMPING POPULATION TRANSFER AND HOLE FILLING SPECTRO-SCOPIES: NEW PROBES OF THE ENERGETIC THRESHOLDS TO CONFORMATIONAL ISOMERIZA-TION II. RESULTS ON ISOLATED, FLEXIBLE BIOMOLECULES

\underline{JASPER R. CLARKSON}, BRIAN C. DIAN, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

Stimulated emission pumping (SEP)-hole filling spectroscopy and SEP-induced population transfer spectroscopy (PTS) have been used to place narrow bounds on the energy thresholds for isomerization between individual reactant-product isomer pairs. SEP-PTS is used to place upper and lower bounds to the barrier heights for various isomer pairs. The thresholds for X→Y and Y→X isomerization are combined to determine the relative energies of the X and Y minima. The combined data from all X→Y isomerizations identify important isomerization pathways on the potential energy surface linking the minima. Application of the aforementioned techniques to several tryptophan analogues, including tryptamine (3-indole ethylamine), 3-indole propionic acid, and melatonin (5-methoxy tryptamine) are presented. Comparison of experimental data to ab initio computed barriers will be made, and the use of the method to test RRKM rate theory applied to conformational isomerization will be discussed.